OpenStudy (konradzuse):

integral of 7/(8x + 1) from 0 to 6

OpenStudy (konradzuse):

\[\int\limits_{0}^{6} \frac{7}{8x+1}\]

OpenStudy (konradzuse):

u = 8x + 1 du = 8dx 1/8 du = dx \[7/8 \int\limits_{0}^{6} \frac{du}{u}\]

OpenStudy (konradzuse):

\[\ln|u| \] for 0 to 6

OpenStudy (konradzuse):

now the issue I'm having is that 7/8[ln(6) - ln(0)] should be infinity, however when I plug it into wolfman It gives me a different answer... http://www.wolframalpha.com/input/?i=integral+of+7%2F%288x+%2B+1%29+from+0+to+6

OpenStudy (konradzuse):

what am i doing wrong?

OpenStudy (anonymous):

all you did wrong is to forget to change your limits of integration... it's not from 0 to 6 anymore

OpenStudy (konradzuse):

oh yeah I forgot, I usually turn it back into x :P.

OpenStudy (anonymous):

:)

OpenStudy (konradzuse):

(ln((8 * 6) + 1) - (ln(1)) which is log(49)

OpenStudy (konradzuse):

* 7/8 = 7log(49)/8

OpenStudy (anonymous):

or 2ln(7)

OpenStudy (konradzuse):

stupid mistake thanks :).

OpenStudy (anonymous):

yw...:)